### v1

@@ A.4 SD-truthfulness for the Three-signal Setting
[…]
@@
Next, we show that if signals are not uniformly self-predicting, the enforced agreement mechanism is not SD-truthful even when the enforcement rule is minimal. We can easily find a counterexample by comparing Eq. 7 with Eq. 8. Suppose W.L.O.G. that $\vartheta^{\tau}_{0,2}<\vartheta^{\tau}_{1,2}$, then Alice can set $t_{1}=t_{2}=t_{3}=0$ and $t_{4}=\max(n_{1},m_{0}-n_{0})$ in which case she can move $t_{4}$ samples from $\mathrm{Bernoulli}(p_{0,2})$ to $\mathrm{Bernoulli}(p_{1,2})$, resulting a score distribution that dominates truth-telling. This completes the proof of necessity. ∎
A.5 Truthfulness of EA
$\max_{\vartheta}\quad$ $\sum_{i,j\in\Sigma}\vartheta_{i,j}\;p_{ij}$
$\vartheta_{i,j}\in\mathbb{Z}_{\geq 0},\quad\forall i,j\in\Sigma.$
Proposition A.1.
A.6 Optimal Enforcement
Proof of Lemma 6.3.
@@ A.6 Optimal Enforcement
Note that for any question, the probability of Alice observing 0 while exerting effort is the marginal distribution $\Pr(X_{a}=0)$, which is identical to the probability of observing 0 while exerting no effort. This implies that under our definition of effort, the probability of Alice observing 0 on $m_{0}$ questions does not depend on the effort. Let $\rho_{0}=\sum_{j\in[n]}(1-X_{a,j})$ be the number of questions where Alice observes 00. The probability that Alice observes 0 on $m_{0}$ questions is
@@
Now, we show that the standard deviation of $S^{EA}(n_{0},e)$ is independent of $n_{0}$. For simplicity, let $\eta_{ij}=e\;p_{ij}+(1-e)\;M^{b}_{j}$. Note that $\eta_{i0}+\eta_{i1}=e\;(p_{i0}+p_{i1})+(1-e)\;(M^{b}_{0}+M^{b}_{1})=1$. Furthermore, the variance of a binomial with parameters $n,p$ is $np(1-p)$. When $m_{0}\leq n_{0}$, the variance of the final score is
@@
Proof of 6.4.
By Lemma 6.3, the standard deviation of the EA mechanism is independent of the enforcement $\Phi$. Therefore, the enforcement that maximizes the sensitivity is the one that maximizes the derivative of the expected score w.r.t. effort $e$. By Eq. 9, the score given by EA can be expressed as the average of many binomial variables. Note that the expectation of a binomial variable with parameter $(n,p)$ is $n\cdot p$. We thus have
@@
[…]
Appendix B Matching Agreement (MA)
The matching agreement mechanism is a follow-up of the correlated agreement mechanism aiming to address more complex cheating strategies (Zhang and Schoenebeck, 2023b). Similar to CA, which determines “agreement” using the delta matrix (see Section 4.3), MA determines “agreement” using a different method depending on the information structure.
Definition 7.
Let $\Gamma$ be a $|\Sigma|\times|\Sigma|\times|\Sigma|$ tensor where
Next, the mechanism randomly samples a bonus question $j$ and a penalty question $k$ uniformly at random. An individual score computed using these two questions is thus $S^{MA}_{i}=T_{\Gamma}(\hat{X}_{a,j},\hat{X}_{b,j},\hat{X}_{b,k})$. In words, the matching agreement mechanism awards Alice a score of 1 if, when predicting her response on question $b$, the posterior conditioned on Bob’s response to the same question provides a better prediction than the posterior conditioned on Bob’s response on a different question $q$. Similar to CA, each bonus question can be paired with $n-1$ penalty questions. Therefore, the final score of MA is the average of $n\cdot(n-1)$ individual scores.
The Sensitivity of MA Under the Partition Rounding Reduction
We further investigate the sensitivity of the MA mechanism under the partition reduction. The MA mechanism requires two questions to compute an individual score, implying that $K=n/2$. Clearly, $S_{\inf}=0$ and $S_{\sup}=1$ by definition. By Eq. 2, to compute the sensitivity, we only need to compute the expected individual score $m_{i}(e)$ and its derivative. When Alice exerts full effort, the probability of obtaining a score of 1 can be computed as follows:
where $(x)^{+}=x$ if $x>0$ and 0 otherwise.
When Alice exerts no effort,
$=\frac{1}{2}$
When signals are binary, we find that MA reduces to CA.
Lemma B.1.
Proof.
$=\sum_{\sigma_{1},\sigma_{2}}(\Delta_{\sigma_{1},\sigma_{2}})^{+}$
An immediate consequence of Lemma B.1 is that the expected score under MA is a linear transformation of the expected score under CA. This further implies that the partition rounding reductions of MA and CA has the same sensitivity under the binary-signal setting.
Proposition B.2.
Proof.
By Lemma B.1, we know that
$\nabla m_{i}^{MA}(e)=\frac{1}{2}\nabla m_{i}^{CA}(e).$
By Eq. 2,
$=\delta^{CA}(e).$
Appendix C Experiments (Continue)
C.1 Varying the Signal Prior
In Fig. 1, we visualize the sensitivities while varying the marginal distribution of the signal prior $\Pr[X_{a}=0]\in[0.1,0.9]$. To do so, we consider the following joint distribution:
$J=\begin{pmatrix}\Pr[X_{a}=0]\cdot 0.75&\Pr[X_{a}=0]\cdot 0.25\\ (1-\Pr[X_{a}=0])\cdot 0.25&(1-\Pr[X_{a}=0])\cdot 0.75\end{pmatrix}.$
In the figure, we vary $\Pr[X_{a}=0]$ with a step size 0.01 from 0.1 to 0.9.
C.2 Information Structure
For Section 7.2 and 7.3, we consider two real-world crowdsourcing datasets, each having a signal space of size $|\Sigma_{1}|=2$ and $|\Sigma_{2}|=4$. The information structure between a pair of agents can be characterized by a Dawid-Skene model (Dawid and Skene, 1979). In particular, each task has an underlying ground truth $\omega$ that is i.i.d. sampled from a prior distribution $W$. Conditioning on $\omega$, each agent receives a signal $\sigma$ according to conditioned distribution $\Gamma$, where $\Gamma_{i,j}=\Pr[\sigma=j\mid\omega=i]$.
In the first dataset (Baba et al., 2018), $18$ agents each labels whether a compound is appropriate or inappropriate to be synthesized. The prior of the ground truth and the conditioned signal distribution are
In the second dataset (Venanzi et al., 2015), $110$ agents provide the annotations of the sentiment of $300$ tweets. There are four ground truth labels and four possible signals where the information structure is

### v2

@@ F.5 EA in the Three-signal Setting
Our goal is to construct a coupling of these Bernoulli variables such that, for any given $\Phi_{X_{a}}$, $\Phi$, and $t_{1},\ldots,t_{4}$, each variable under truth-telling with manipulation matrix $\vartheta^{\tau}(\rho^{*})$ has a weakly higher probability of success than its coupled counterpart under $\vartheta$. We denote this coupling by $Z^{\tau}_{i,j}\xrightarrow{k}Z^{\theta}_{i^{\prime},j^{\prime}}$ to indicate that we pair $k$ Bernoulli variables associated with $\vartheta^{\tau}(\rho^{*})$—each drawn from $\mathrm{Bernoulli}(p_{ij})$—with $k$ Bernoulli variables associated with $\vartheta$—each drawn from $\mathrm{Bernoulli}(p_{i^{\prime}j^{\prime}})$. Our coupling is:
@@
Next, we show that if signals are not uniformly self-predicting, the enforced agreement mechanism is not SD-truthful even when the enforcement rule is minimal. We can easily find a counterexample by comparing Equation 6 with Equation 7. Suppose W.L.O.G. that $\vartheta^{\tau}_{0,2}<\vartheta^{\tau}_{1,2}$, then Alice can set $t_{1}=t_{2}=t_{3}=0$ and $t_{4}=\max(n_{1},m_{0}-n_{0})$ in which case she can move $t_{4}$ samples from $\mathrm{Bernoulli}(p_{0,2})$ to $\mathrm{Bernoulli}(p_{1,2})$, resulting a score distribution that dominates truth-telling. This completes the proof of necessity. ∎
F.6 The Optimal Enforcement
Proof of Lemma C.5.
@@ F.6 The Optimal Enforcement
Note that for any question, the probability of Alice observing 0 while exerting effort is the marginal distribution $\Pr(X_{a}=0)$, which is identical to the probability of observing 0 while exerting no effort. This implies that under our definition of effort, the probability of Alice observing 0 on $m_{0}$ questions does not depend on the effort. Let $\rho_{0}=\sum_{j\in[n]}(1-X_{a,j})$ be the number of questions where Alice observes 0. The probability that Alice observes 0 on $m_{0}$ questions is
@@
Now, we show that the standard deviation of $S^{EA}(n_{0},e)$ is independent of $n_{0}$. For simplicity, let $\eta_{ij}=e\;p_{ij}+(1-e)\;M^{b}_{j}$. Note that $\eta_{i0}+\eta_{i1}=e\;(p_{i0}+p_{i1})+(1-e)\;(M^{b}_{0}+M^{b}_{1})=1$. Furthermore, the variance of a binomial with parameters $n,p$ is np(1p)np(1-p). When $m_{0}\leq n_{0}$, the variance of the final score is
@@
Proof of C.6.
By Lemma C.5, the standard deviation of the EA mechanism is independent of the enforcement $\Phi$. Therefore, the enforcement that maximizes the sensitivity is the one that maximizes the derivative of the expected score w.r.t. effort $e$. By Equation 8, the score given by EA can be expressed as the average of many binomial variables. Note that the expectation of a binomial variable with parameter $(n,p)$ is $n\cdot p$. We thus have
@@
$\Leftrightarrow\quad$ $F(n_{0})\left(\eta^{\prime}_{10}-\eta^{\prime}_{11}\right)+(1-F(n_{0}))\left(\eta^{\prime}_{00}-\eta^{\prime}_{01}\right)>0$
$\Leftrightarrow\quad$ $F(n_{0})<\frac{\eta^{\prime}_{00}-\eta^{\prime}_{01}}{\eta^{\prime}_{00}-\eta^{\prime}_{01}+\eta^{\prime}_{11}-\eta^{\prime}_{10}}.$